OpenStudy (anonymous):

Find the exact value of the expression without using a calculator. cos(2tan^-1(12/5)) Thanks

OpenStudy (mertsj):



OpenStudy (mertsj):

Are you sure you have posted that problem correctly?

OpenStudy (anonymous):

\[\cos(2\tan^{-1} 12/5)\]

OpenStudy (mertsj):

\[Let \tan^-1(12/5)=x\] cos2x= 2cos^2x-1=2cos^2[tan^-1(12/5)]-1=2(5/13)^2-1=50/169-1=-119/169